OpenStudy (anonymous):

given the acceleration, initial velocity, and initial position of a body moving along a coordinate line at time t, find the body's position at time t.

OpenStudy (anonymous):

a=12, v (0)=7, s (0)=5

OpenStudy (anonymous):

*s (0)=-1

OpenStudy (owlcoffee):

Okay, this is pretty much physics and also a little bit of thinking. How would you use the data they gave to find the time?

OpenStudy (anonymous):

I'm not sure. What's the first step of doing this problem?

OpenStudy (anonymous):

First I would get the antiderivative correct?

OpenStudy (owlcoffee):

Correct, in order to find the movement.

OpenStudy (anonymous):

V=9.8t+9 s=4.9^2+9t+(-1)

OpenStudy (anonymous):

What up dude I really need some help

OpenStudy (anonymous):

V=9.8t+9 s=4.9^2+9t+(-1)